OpenStudy (anonymous):

WILL AWARD METAL!! Convert the polar representation of this complex number into its rectangular form Z=4(cos 150 degrees+isin 150 degrees) a)-2(sqrt 3),2 b) -2, 2(sqrt 3) c) -4, 4(sqrt 3) d) -2(sqrt 3), -2 Please and thank you :)

OpenStudy (anonymous):

@Thesn1 Please help?

OpenStudy (anonymous):

Its A!

OpenStudy (anonymous):

see with me my friend next time i ll get angry if you dont solve those kind of question alone step 1 : 150deg=180deg+(-30deg) step 2 we replace cos150 by -cos (30) and sin150 by sin30 then the write answer is a

OpenStudy (anonymous):

Could you help me with one more? It is write the product in standard form (1+3i)(5-i)? I guess I don't know what the standard form is.

OpenStudy (anonymous):

(1+3i)(5-i)=5-i+15i+3=8+14i

OpenStudy (anonymous):

i have to go now i ll leave you a contact in message for more help

OpenStudy (anonymous):

Okay thank you!